the moment of inertia of a solid flywheel of radius 0.1m about its axis is 0.1kgm2. a tangential force of 2 kg wt is applied through the circumference of flywheel with the help of string and mass M. the acceleration of the mass is

Grade:11

1 Answers

Adikesh Kumar
7 years ago
ANS: 2m/s2
equate the torque 
I alpha = force * r  (r = radius of wheel)
0.1*a/r = 2*10* r    ( a = accelaration)
therefore a=2m/s ( r = 0.1m)